# D. Leykekhman, B. Vexler and J. Wagner

### v1

@@ 6 Full Discretization of the Optimal Control Problem
$\left({\bf{q}}-\pi_{d}{\bf{q}},\mathbf{r}\right)_{I\times\Omega}=0\qquad\forall\mathbf{r}\in\mathbf{Q}_{0}.$ (6.2)
@@
Minimize $\text{Minimize }\;J({\bf{q}}_{\sigma},{\bf{u}}_{\sigma})=\frac{1}{2}\lVert{\bf{u}}_{\sigma}-{\bf{u}}_{d}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+\frac{\alpha}{2}\lVert{\bf{q}}_{\sigma}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}$ (6.4a)
@@
$G_{\bf{w}}({\bf{u}}_{kh}(\pi_{d}\tilde{\bf{q}}))=G_{\bf{w}}({\bf{u}}_{kh}(\pi_{d}\tilde{\bf{q}}-\tilde{\bf{q}}))+G_{\bf{w}}({\bf{u}}_{kh}(\tilde{\bf{q}}))<C\|\pi_{d}\tilde{\bf{q}}-\tilde{\bf{q}}\|_{L^{2}(I;L^{2}(\Omega))}+\beta-\delta.$
As 3 guarantees $k\to 0$ as $h\to 0$, it holds $\|\pi_{d}\tilde{\bf{q}}-\tilde{\bf{q}}\|_{L^{2}(I;L^{2}(\Omega))}\to 0$ for $h\to 0$. This implies that for $h$ small enough, we have $C\|\pi_{d}\tilde{\bf{q}}-\tilde{\bf{q}}\|_{L^{2}(I;L^{2}(\Omega))}\leq\delta$ and as a consequence $G_{\bf{w}}({\bf{u}}_{kh}(\pi_{d}\tilde{\bf{q}}))<\beta$. ∎
@@
$\|\bar{{\bf{q}}}_{\sigma}\|_{L^{\infty}(I;L^{2}(\Omega))}+\|\bar{{\bf{u}}}_{\sigma}\|_{L^{2}(I;L^{2}(\Omega))}+\|\bar{\mu}_{\sigma}\|_{(C(\bar{I}))^{*}}\leq C.$
@@
$J(\bar{{\bf{q}}}_{\sigma},\bar{{\bf{u}}}_{\sigma})$ $\leq J(\pi_{d}\tilde{{\bf{q}}},\bar{{\bf{u}}}_{kh}(\pi_{d}\tilde{{\bf{q}}}))$
$=\frac{1}{2}\lVert\bar{{\bf{u}}}_{kh}(\pi_{d}\tilde{{\bf{q}}})-{\bf{u}}_{d}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+\frac{\alpha}{2}\lVert\pi_{d}\tilde{{\bf{q}}}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}$
$\leq\lVert\bar{{\bf{u}}}_{kh}(\tilde{{\bf{q}}})-{\bf{u}}(\tilde{{\bf{q}}})\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+\lVert{\bf{u}}(\tilde{{\bf{q}}})-{\bf{u}}_{d}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+\frac{\alpha}{2}\lVert\tilde{{\bf{q}}}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}$
$\leq C(T,\tilde{{\bf{q}}},{\bf{u}}_{d}).$
As a result we obtain $\|\bar{{\bf{q}}}_{\sigma}\|_{L^{2}(I;L^{2}(\Omega))}+\|\bar{{\bf{u}}}_{\sigma}\|_{L^{2}(I;L^{2}(\Omega))}\leq C.$ The proof of $\|\bar{\mu}_{\sigma}\|_{C(\bar{I})^{*})}\leq C$ and $\|\bar{\bf{q}}_{\sigma}\|_{L^{\infty}(I;L^{2}(\Omega))}\leq C$ is then accomplished by following the same steps as the proof of Lemma 5.7, making use of $G_{\bf{w}}({\bf{u}}_{kh}({\pi_{d}{\bf{p}}}))\leq\beta-\frac{1}{4}\delta$ for $h$ sufficiently small.
@@
$\alpha\lVert\bar{{\bf{q}}}_{kh}-\bar{{\bf{q}}}_{\sigma}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+\lVert\bar{{\bf{u}}}_{kh}-\bar{{\bf{u}}}_{\sigma}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}\leq Ch^{2}.$
@@
$\alpha\|\bar{{\bf{q}}}_{kh}-\bar{{\bf{q}}}_{\sigma}\|^{2}_{L^{2}(I;L^{2}(\Omega))}\leq(\bar{{\bf{z}}}_{kh}-\bar{{\bf{z}}}_{\sigma},\bar{{\bf{q}}}_{\sigma}-\bar{{\bf{q}}}_{kh})_{I\times\Omega}+(\alpha\bar{{\bf{q}}}_{\sigma}+\bar{{\bf{z}}}_{\sigma},\pi_{d}\bar{{\bf{q}}}_{kh}-\bar{{\bf{q}}}_{kh})_{I\times\Omega}:=I_{1}+I_{2}.$ (6.9)
@@
$\leq-\lVert\bar{{\bf{u}}}_{kh}-\bar{{\bf{u}}}_{\sigma}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+\langle\beta-G_{\bf w}(\bar{{\bf{u}}}_{kh}),\bar{\mu}_{kh}\rangle+\langle\beta-G_{\bf w}(\bar{{\bf{u}}}_{\sigma}),\bar{\mu}_{\sigma}\rangle$
$=-\lVert\bar{{\bf{u}}}_{kh}-\bar{{\bf{u}}}_{\sigma}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}.$
@@
$\leq Ch^{2}\|\nabla\bar{\bf{z}}_{\sigma}\|_{L^{2}(I;L^{2}(\Omega))}\|\nabla\bar{\bf{q}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}.$
Using that $\|{\rm NewA}\bar{{\bf{q}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}\leq\alpha^{-1}\|{\rm NewA}\bar{{\bf{z}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}$, the stability of $P_{ad}$ and the stability of solutions to the fully discrete dual problem, pointed out in Remark 4.3, gives
$I_{2}\leq Ch^{2}\left(\|\bar{{\bf{u}}}_{\sigma}-\bar{{\bf{u}}}_{d}\|_{L^{2}(I;L^{2}(\Omega))}+\|\bar{\mu}_{\sigma}\|_{L^{1}(I)}\|{\bf{w}}\|_{L^{2}(\Omega)}\right)\left(\|\bar{{\bf{u}}}_{kh}-\bar{{\bf{u}}}_{d}\|_{L^{2}(I;L^{2}(\Omega))}+\|\bar{\mu}_{kh}\|_{L^{1}(I)}\|{\bf{w}}\|_{L^{2}(\Omega)}\right).$
Now the boundedness of $\|\bar{{\bf{u}}}_{\sigma}\|_{L^{2}(I;L^{2}(\Omega))}$, $\|\bar{{\bf{u}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}$, $\|\bar{\mu}_{\sigma}\|_{L^{1}(I)}$, and $\|\bar{\mu}_{kh}\|_{L^{1}(I)}$ from Lemmas 5.7 and 6.4 finish the proof. ∎
@@
$\sqrt{\alpha}\lVert\bar{{\bf{q}}}-\bar{{\bf{q}}}_{\sigma}\rVert_{L^{2}(I;L^{2}(\Omega))}+\lVert\bar{{\bf{u}}}-\bar{{\bf{u}}}_{\sigma}\rVert_{L^{2}(I;L^{2}(\Omega))}\leq C\,\ell_{k}(k^{\frac{1}{2}}+h),\quad\ell_{k}=\ln{\frac{T}{k}}.$
@@ 8 Numerical results
In the following, we present two numerical examples, studying the orders of convergence for the optimal control problem. We first present an example for smooth data. These results can be compared to the numerical example in [29], where the weight function $w(x_{1},x_{2})=\sin{(\pi x_{1})}\sin{(\pi x_{2})}$ on the unit square was used. As this weight is in $H^{1}_{0}(\Omega)\cap H^{2}(\Omega)$ almost first order convergence in time was observed, which we again can see for the Stokes optimal control problem. In the second example, we study an example where less regularity of the data is available, leading to a reduced order of convergence.

### v2

@@ 6 Full Discretization of the Optimal Control Problem
$\left({\bf{q}}-\pi_{d}{\bf{q}},\mathbf{r}\right)_{I\times\Omega}=0\qquad\forall\mathbf{r}\in\mathbf{Q}_{0},$ (6.2)
which by definition is stable in ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}$, i.e., satisfies
$\|\pi_{d}{\bf{q}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\leq\|{\bf{q}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}.$ (6.3)
@@
Minimize $\text{Minimize }\;J({\bf{q}}_{\sigma},{\bf{u}}_{\sigma})=\frac{1}{2}\lVert{\bf{u}}_{\sigma}-{\bf{u}}_{d}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\frac{\alpha}{2}\lVert{\bf{q}}_{\sigma}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$ (6.5a)
@@
$G_{\bf{w}}({\bf{u}}_{kh}(\pi_{d}\tilde{\bf{q}}))=G_{\bf{w}}({\bf{u}}_{kh}(\pi_{d}\tilde{\bf{q}}-\tilde{\bf{q}}))+G_{\bf{w}}({\bf{u}}_{kh}(\tilde{\bf{q}}))<C\|\pi_{d}\tilde{\bf{q}}-\tilde{\bf{q}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\beta-\delta.$
As 3 guarantees $k\to 0$ as $h\to 0$, it holds $\|\pi_{d}\tilde{\bf{q}}-\tilde{\bf{q}}\|_{L^{2}(I;L^{2}(\Omega))}\to 0$ for $h\to 0$. This implies that for $h$ small enough, we have $C\|\pi_{d}\tilde{\bf{q}}-\tilde{\bf{q}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\leq\delta$ and as a consequence $G_{\bf{w}}({\bf{u}}_{kh}(\pi_{d}\tilde{\bf{q}}))<\beta$. ∎
@@
$\|\bar{{\bf{q}}}_{\sigma}\|_{L^{\infty}(I;L^{2}(\Omega))}+\|\bar{{\bf{u}}}_{\sigma}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\|\bar{\mu}_{\sigma}\|_{(C(\bar{I}))^{*}}\leq C.$
@@
$J(\bar{{\bf{q}}}_{\sigma},\bar{{\bf{u}}}_{\sigma})$ $\leq J(\pi_{d}\tilde{{\bf{q}}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf{u}}_{kh}}(\pi_{d}\tilde{{\bf{q}}}))$
$=\frac{1}{2}\lVert{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf{u}}_{kh}}(\pi_{d}\tilde{{\bf{q}}})-{\bf{u}}_{d}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\frac{\alpha}{2}\lVert\pi_{d}\tilde{{\bf{q}}}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$
$\leq\lVert{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf{u}}_{kh}}(\pi_{d}\tilde{{\bf{q}}})\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\lVert{\bf{u}}_{d}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\frac{\alpha}{2}\lVert\tilde{{\bf{q}}}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$
$\leq C(T,\tilde{{\bf{q}}},{\bf{u}}_{d}),$
where in the last step, we have used the discrete stability result for ${\bf{u}}_{kh}$ from Theorem 4.2 and the stability of $\pi_{d}$ from (6.3). As a result we obtain $\|\bar{{\bf{q}}}_{\sigma}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\|\bar{{\bf{u}}}_{\sigma}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\leq C.$ The proof of $\|\bar{\mu}_{\sigma}\|_{C(\bar{I})^{*})}\leq C$ and $\|\bar{\bf{q}}_{\sigma}\|_{L^{\infty}(I;L^{2}(\Omega))}\leq C$ is then accomplished by following the same steps as the proof of Lemma 5.7, making use of $G_{\bf{w}}({\bf{u}}_{kh}({\pi_{d}{\bf{p}}}))\leq\beta-\frac{1}{4}\delta$ for $h$ sufficiently small.
@@
$\alpha\lVert\bar{{\bf{q}}}_{kh}-\bar{{\bf{q}}}_{\sigma}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\lVert\bar{{\bf{u}}}_{kh}-\bar{{\bf{u}}}_{\sigma}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\leq Ch^{2}.$
@@
$\alpha\|\bar{{\bf{q}}}_{kh}-\bar{{\bf{q}}}_{\sigma}\|^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\leq(\bar{{\bf{z}}}_{kh}-\bar{{\bf{z}}}_{\sigma},\bar{{\bf{q}}}_{\sigma}-\bar{{\bf{q}}}_{kh})_{I\times\Omega}+(\alpha\bar{{\bf{q}}}_{\sigma}+\bar{{\bf{z}}}_{\sigma},\pi_{d}\bar{{\bf{q}}}_{kh}-\bar{{\bf{q}}}_{kh})_{I\times\Omega}:=I_{1}+I_{2}.$ (6.10)
@@
$\leq-\lVert\bar{{\bf{u}}}_{kh}-\bar{{\bf{u}}}_{\sigma}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\langle\beta-G_{\bf w}(\bar{{\bf{u}}}_{kh}),\bar{\mu}_{kh}\rangle+\langle\beta-G_{\bf w}(\bar{{\bf{u}}}_{\sigma}),\bar{\mu}_{\sigma}\rangle$
$=-\lVert\bar{{\bf{u}}}_{kh}-\bar{{\bf{u}}}_{\sigma}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}.$
@@
$\leq Ch^{2}\|\nabla\bar{\bf{z}}_{\sigma}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\|\nabla\bar{\bf{q}}_{kh}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}.$
Using that $\|{\rm NewA}\bar{{\bf{q}}}_{kh}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\leq\alpha^{-1}\|{\rm NewA}\bar{{\bf{z}}}_{kh}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$, which holds due to the projection formula (5.5d) and the stability of $P_{ad}$ in $H^{1}(\Omega)$, see [27, Theorem A.1], [2, Theorem 5.8.2], and the stability of solutions to the fully discrete dual problem, pointed out in Remark 4.3, gives
$I_{2}\leq Ch^{2}\left(\|\bar{{\bf{u}}}_{\sigma}-\bar{{\bf{u}}}_{d}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\|\bar{\mu}_{\sigma}\|_{L^{1}(I)}\|{\bf{w}}\|_{L^{2}(\Omega)}\right)\left(\|\bar{{\bf{u}}}_{kh}-\bar{{\bf{u}}}_{d}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\|\bar{\mu}_{kh}\|_{L^{1}(I)}\|{\bf{w}}\|_{L^{2}(\Omega)}\right).$
Now the boundedness of $\|\bar{{\bf{u}}}_{\sigma}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$, $\|\bar{{\bf{u}}}_{kh}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$, $\|\bar{\mu}_{\sigma}\|_{L^{1}(I)}$, and $\|\bar{\mu}_{kh}\|_{L^{1}(I)}$ from Lemmas 5.7 and 6.4 finish the proof. ∎
@@
$\sqrt{\alpha}\lVert\bar{{\bf{q}}}-\bar{{\bf{q}}}_{\sigma}\rVert_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\lVert\bar{{\bf{u}}}-\bar{{\bf{u}}}_{\sigma}\rVert_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\leq C\,\ell_{k}(k^{\frac{1}{2}}+h),\quad\ell_{k}=\ln{\frac{T}{k}}.$
@@ 8 Numerical results
In the following, we present three numerical examples, studying the orders of convergence for the optimal control problem. We first present an example for smooth data. These results can be compared to the numerical example in [33], where the weight function $w(x_{1},x_{2})=\sin{(\pi x_{1})}\sin{(\pi x_{2})}$ on the unit square was used. As this weight is in $H^{1}_{0}(\Omega)\cap H^{2}(\Omega)$ almost first order convergence in time was observed, which we again can see for the Stokes optimal control problem. In the second example, we study an example where less regularity of the data is available, leading to a reduced order of convergence. Lastly, we consider an example with simultaneous control and state constraints. Throughout this section, we will analyze the errors of the full discretization $\bar{\bf{q}}-\bar{\bf{q}}_{\sigma}$. For details regarding the implementation of a variational discretization in the presence of control constraints, we refer to [21], [22, Chapter 3.2.5] and the references therein.